The (partial) cost sheet for the single product manufactured at Vienna Company follows.

Direct labor (25 hours @ $25) $ 625
Variable overhead (25 hours @ $2) 50
Fixed overhead (25 hours @ $4) 100


The master budget level of production is 98,000 direct-labor hours, which is also the production volume used to compute the fixed overhead application rate. Other information available for operations over the past accounting period include the following.

Actual variable overhead incurred $ 148,000
Actual fixed overhead incurred 408,800
Direct labor efficiency variance 178,000 U
Variable overhead price variance 46,000 F

Required:
a. What was the variable overhead efficiency variance?
b. What was the fixed overhead price variance?
c. What was the fixed overhead production volume variance?

(For all requirements, indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)


a. Variable overhead efficiency variance $ U
b. Fixed overhead price variance U
c. Fixed overhead production volume variance U
Answer #1

a. Variable overhead price variance = (Actual hours* standard variable overhead rate per hour) - Actual variable overhead

46000 = (Actual hours x 2) - 148000

Actual hours = 97000

Variable Overhead Efficiency Variance = Standard overhead rate x (Actual hours - Standard hours)

= (98000-97000)*2 = $2000 Favorable

b. Fixed overhead price variance = Actual fixed overhead incurred - budgeted fixed overhead

= 408800 - (98000*4)

= $17800 Unfavorable

c. Fixed overhead production volume variance = (Actual hours - budgeted hours)*Budgeted rate

= (97000-98000)*4

= $4000 Favorable